OpenStudy (anonymous):

Did someone make Supplementary Problems, in particular exercise 1A-4 (vectors and matrices, MIT 18.02 course)? If yes, why OX = s*OP+r*OQ???!!!

OpenStudy (phi):

Perhaps a concrete example will help. let r= 1/5 and s= 4/5 and r+s=1